jennylove:

A sonar pulse is sent to the ocean bottom at 25°C, and it takes 1.20 s for the echo to come back. How deep is the bottom? 1.79 km 0.92 km 1.84 km 0.90 km Speed of Sound Waves in Various Media Medium m/s Air (0°C) 331 Air (20°C) 343 Helium (0°C) 972 Water (25°C) 1,493 Seawater (25°C) 1,533 Copper (25°C) 3,560 Iron (25°C) 5,130

jennylove:

i think its 0.92

justjm:

I actually got C - 1.84 km. Did you use Seawater?

justjm:

oh shoot wait a sec

justjm:

you are correct

justjm:

forgot to divide by 2 after bc it comes back lol

jennylove:

I did Distance = Velocity * Time = 1560 * 1.20 meter = 1872 meter The depth of the bottom = 1872/2 meter = 936 meter = 0.92 km

jennylove:

yes lol

jennylove:

awesome! glad we figured that out(:

justjm:

Yep np